Discuss in detail all cases of the roots of a second order linear differential equation with constant

coefficients.


1
Expert's answer
2021-07-13T08:56:29-0400

Given the second order linear homogeneous equations with constant coefficients


"ay''+by'+c=0, a\not=0"

Let the characteristic equation of the differential equation be


"ar^2+br+c=0"

Case 1 Two distinct real roots

When "b^2-4ac>0," the characteristic polynomial has two distinct real roots "r_1, r_2." They give two distinct solutions "y_1=e^{r_1t}" and "y_2=e^{r_2t}."

Therefore, a general solution of the second order linear homogeneous equations with constant coefficients is


"y(t)=c_1e^{r_1t}+c_2e^{r_2t}"


Case 2 Two complex conjugate roots

When "b^2-4ac<0," the characteristic polynomial has two complex roots, which are conjugates, "r_1=\lambda+\mu i, r_2=\lambda-\mu i" ("\lambda,\mu" are real numbers, "\mu>0" ).

As before they give two linearly independent solutions "y_1=e^{r_1t}" and "y_2=e^{r_2t}." Consequently the linear combination "y(t)=c_1e^{r_1t}+c_2e^{r_2t}" will be a general solution. 

Using the Euler's Formula we have


"y(t)=c_3e^{\lambda t}\cos (\mu t)+c_4e^{\lambda t}\sin (\mu t)"

Case 3 One repeated real root

When "b^2-4ac=0," the characteristic polynomial has a single repeated real root "r=-\dfrac{b}{2a}." The general solution in the case of a repeated real root "r" is


"y(t)=c_1e^{rt}+c_2te^{rt}."